# Benchmarking the Text-to-SQL Capability of Large Language Models: A Comprehensive Evaluation

### v1

@@ Abstract
Large Language Models (LLMs) have emerged as a powerful tool in advancing the Text-to-SQL task, significantly outperforming traditional methods. Nevertheless, as a nascent research field, there is still no consensus on the optimal prompt templates and design frameworks. Additionally, existing benchmarks inadequately explore the performance of LLMs across the various sub-tasks of the Text-to-SQL process, which hinders the assessment of LLMs’ cognitive capabilities and the optimization of LLM-based solutions. To address the aforementioned issues, we firstly construct a new dataset designed to mitigate the risk of overfitting in LLMs. Then we formulate five evaluation tasks to comprehensively assess the performance of diverse methods across various LLMs throughout the Text-to-SQL process. Our study highlights the performance disparities among LLMs and proposes optimal in-context learning solutions tailored to each task. These findings offer valuable insights for enhancing the development of LLM-based Text-to-SQL systems.
@@ 1 Introduction
Text-to-SQL, which involves the automatic transformation of natural language (NL) questions into structured SQL statements, is a pivotal component in facilitating seamless user interaction with databases [35]. Previous approaches to this task primarily focus on pattern matching between natural language and SQL statements, utilizing machine learning models to acquire the mapping between the two [60, 21]. However, the introduction and rapid advancement of Large Language Models (LLMs) have brought about a substantial transformation in this field [9, 33]. LLMs have emerged as powerful tools [38, 18, 43], showcasing tremendous potential in comprehending complex NL questions and generating accurate SQL statements. By combining advanced reasoning techniques and in-context learning capabilities, LLMs have significantly pushed the boundaries of the state-of-the-art in this domain, outperforming traditional methods by a considerable margin.
Despite the continuous improvement of LLM-based methods in various benchmarks such as Spider [56] and BIRD [24], there remains a critical gap in the systematic benchmarking of these solutions [17, 35, 19]. The absence of a comprehensive framework for evaluating LLMs in Text-to-SQL complicates the design and assessment of effective systems. The risk of overfitting, particularly for LLMs trained on coding tasks and open-source datasets, poses a significant challenge to the reliability of benchmark evaluations [37]. Additionally, the optimal prompt engineering strategies that play a crucial role in guiding LLMs to generate accurate SQL queries are yet to be determined. Although various design schemes are explored in different methods [13], there is no consensus on the most effective prompt template. Furthermore, the current benchmarks, while comprehensive in their assessment of end-to-end Text-to-SQL tasks, have not yet provided a detailed exploration of the models’ performance across the various sub-tasks and components of the Text-to-SQL process [32, 27]. A detailed exploration of these sub-tasks is crucial for a thorough evaluation of LLMs’ cognitive capabilities and their role in facilitating the Text-to-SQL process. Therefore, it is necessary to develop a more granular benchmarking approach that can accurately reflect the multifaceted nature of Text-to-SQL and inform the creation of more effective LLM-based solutions.
@@ 2.1 Traditional Learning-based Text-to-SQL Methods
Numerous traditional learning-based Text-to-SQL methods existed before the emergence of LLMs [60, 52, 29]. These methods can be essentially divided into non-seq2seq and seq2seq methods according to their network architecture. In Non-seq2seq methods, representative works [46, 2, 1, 15] typically employed a relation-aware self-attention mechanism as encoder to learn representations of questions and schemas, and then used a grammar-based decoder to generate the SQL as an abstract syntax tree [54], or utilized a sketch-based decoder to obtain the SQL via slot-filling [8, 16, 5]. These methods can further benefit from leveraging pre-trained language models like BERT and its extensions [7, 28, 55] for input embedding initialization. As another line of research, seq2seq methods [36, 26, 34, 23] directly translate NL questions into SQL queries through transformer-based [45] architectures in an end-to-end manner. These methods obtained competitive performance by fine-tuning rather than training from scratch, with less effort. They typically used smarter decoding techniques (e.g., constraining the predictions of the decoder [39], or schema-aware denoising [50]) to prevent the production of invalid SQLs. RESDSQL [21] further decoupled the intertwined schema linking (determining the schema items like tables and columns in a SQL) and the skeleton parsing (determining the SQL keywords) process, which alleviated the difficulty of Text-to-SQL. Apart from network architectures, the paradigms employed by these methods varied greatly in terms of input encoding, output decoding, neural training, output refinement, making Text-to-SQL a flourishing research area [17]. However, the highest accuracy on the Spider leaderboard achieved by traditional learning-based methods is 79.9%, which is still far from being a reliable Text-to-SQL parser.
@@ 2.2 LLM-based Text-to-SQL Methods
With the advancements in LLMs, researchers are increasingly interested in creating a natural language interface for relational databases through the powerful linguistic and coding capabilities of LLMs, forge a new trend of LLM-based Text-to-SQL. Given the Powerful zero-shot reasoning and domain generalization abilities of LLMs, these methods successively refreshed the record on the cross-domain Spider leaderboard. C3 [9], a zero-shot Text-to-SQL method based on ChatGPT, provided treatment in terms of model input, bias and output, reaching an 82.3% execution accuracy on Spider leaderboard. DIN-SQL [33] proposed decomposing the Text-to-SQL task into smaller sub-tasks effectively and achieved an accuracy of 85.3% on Spider. DAIL-SQL [13] refreshes the accuracy of Spider with 86.6% through both supervised fine-tuning and a systematic study of in-context learning. It explored how to select the most helpful example and organize them properly in prompts in a few-shot scenario. Similarly, other studies investigated the selection of few-shot demonstrations by synthesising in-domain examples [3] and retrieving question skeletons [14]. MAC-SQL [47] utilized multi-agent collaboration for Text-to-SQL tasks and reached an accuracy of 59.6% on more challenging BIRD.
Several other studies focused on special yet non-trivial scenarios, aiming to expand the scope of the Text-to-SQL task. DBCopilot [48] considered a more realistic problem, wherein it dealt with large-scale schemas, characterized by massive databases and a large number of tables. It proposed the use of a lightweight seq2seq copilot model for schema-routing, increasing scalability in comparison to traditional schema-linking. ACT-SQL [58] designed a chain-of-thought (CoT) [49] prompt to improve the reasoing ability when generating SQLs, and extended to the multi-turn Text-to-SQL task [22].
In summary, these methods primarily focused on improving the performance of the overall Text-to-SQL task through several aspects of the sub-tasks of interest in this paper. However, these methods 1) did not individually and systematically study the performance of sub-tasks, and 2) typically relied on experiments using only OpenAI LLMs (gpt-3.5-turbo, gpt4, etc.), without verifying the robustness of their approaches when applied to other open-source LLMs.
@@ 2.3 Datasets and Evaluation Metrics
WikiSQL [60] is considered the first large-scale dataset enabling the training and evaluation of learning-based Text-to-SQL methods, as well as offering a standardized benchmark for straightforward comparison across various methods. It is also known as a cross-domain dataset, featuring over 25,000 tables and 80,000 question-SQL pairs comprising a variety of domains derived from Wikipedia.
@@
Subsequently, most of the recent Text-to-SQL works are done on Spider [56] because it is widely acknowledged as the most challenging cross-domain benchmark. It comprises 10,181 queries covering 138 different domains, involves multi-table queries (embodied by JOIN), complex SQL clauses (ORDER BY, GROUP BY, and HAVING, etc.) and nested SQLs. Several variants of Spider have been designed to evaluate the adaptability of Text-to-SQL methods [40, 12]. Spider-Realistic [6] removes the explicit mention of column names in the NL questions while keeping the SQL queries unchanged, which is more aligned with the use-case of Text-to-SQL in practice. Spider-Syn [10] replaces some schema words in the NL question with their synonyms that reflect real-world question paraphrases. Eliminating such explicit correspondence between NL questions and table schemas poses a greater challenge of Text-to-SQL methods. Spider-DK [11] integrates artificial domain knowledge, aiming to investigate the robustness of text-to-SQL models when the questions require rarely observed domain knowledge.
@@
Given that Spider was curated for benchmarking Text-to-SQL methods while diverging from real-world scenarios, KaggleDBQA [20] constructed a small-scale cross-domain dataset from Kaggle, highlighting the practicality and diversity of actual web-sourced databases. KaggleDBQA firstly comprises documentation and metadata of databases, raising an question of how this extra information could be used to improve the performance.
@@
The newest seminal benchmark is BIRD [24], comprising 12,751 Text-to-SQL pairs and 95 databases with a size of 33.4 GB. Different from previous datasets, it is primarily designed to evaluate 1) the SQL execution efficiency in large-scale databases and 2) the ability of using external knowledge evidence. In this paper, we construct a novel dataset built upon BIRD and use it in the evaluation (see Section 3.3).
@@
Two primary evaluation metrics for assessing the accuracy of SQLs on Spider are Exact Matching(EM) and Execution Accuracy(EX). EM measures whether the predicted query as a whole is equivalent to the gold query. It is possible to encounter false negative evaluations since a question might be solvable by multiple syntactically different but semantically identical SQL statements. EX is a more widely used metric, measures whether the result of executing the predicted query matches the gold value. We use EX to evaluate the accuracy of SQLs in this paper. Additionally, BIRD further proposed Valid Efficiency Score (VES), an integrated metric assessing both accuracy of execution results (i.e., EX) and the execution efficiency of SQL queries. To increase the VES, a Text-to-SQL method requires enhancing execution efficiency of SQL queries without sacrificing EX. We use this metric to assess the SQL optimization in Section 4.3. Refer to Appendix C.1 for detailed definitions of evaluation metrics.
@@ 3.2 Evaluation Models
Our benchmarking study evaluates the performance of two distinct categories of LLMs with varying parameter sizes: general-purpose and coding-specific. General-purpose LLMs are designed for versatile text generation and comprehension across diverse domains, trained on extensive internet text datasets. Specifically, ChatGPT (gpt-35-turbo-16k)  [^1] footnote, LLaMa2-Chat-70B [44], InternLM-70B  [^2] https://internlm.intern-ai.org.cnand InternLM2-20B are selected as the main baseline models. Coding-specific LLMs are fine-tuned and optimized for programming scenarios, excelling in code generation and technical language understanding. In this paper, the performance analysis of Codellama-34B [37] and SQLCoder-34B  [^3] https://github.com/defog-ai/sqlcoder are provided.
@@ 3.3 Dataset Construction
Specifically, our construction process involves a systematic analysis of the original BIRD dataset, identifying queries of varying difficulty levels and involving different numbers of tables (1, 2, 3, and more than 3). We modify and expand these queries by altering table and column names, as well as filtering conditions, to create a more diverse set of challenges. In cases where the original dataset lacks sufficient examples with four or more tables (there are only 20 instances in BIRD-Dev dataset), queries that involved three tables are expanded to four. As shown in Table 2, this process generates 50 new instances for each category, resulting in the “BigTable-0.2k” dataset. Moreover, each instance in the dataset underwent mutual verification by at least two individuals to ensure the accuracy of the data.
@@ 4.1 Text2SQL
Unlike previous learning-based studies, the primary challenge in LLM-based Text-to-SQL is the design of an effective prompt template $\mathcal{P}$ (as introduced in Section 3.1) for LLMs to generate accurate SQL queries, known as prompt engineering. Researchers have evaluated a variety of prompt templates [13]. However, these representations lack uniformity in their structure, making it difficult to find out how a specific feature within a prompt template impact performance.
@@
prompt template SQLCoder-34B InternLM-70B Codellama-34B Llama2-Chat-70B
DDL-HTML-Chat 57.8 65.0 63.7 49.6
DDL-HTML-Complete 61.8 53.8 65.2 50.2
DDL-MD-Chat 63.2 66.3 68.4 48.7
DDL-MD-Complete 62.4 64.1 69.8 46.8
DDL-Coding-Chat 60.3 66.1 67.1 48.4
DDL-Coding-Complete 59.7 62.9 66.9 53.4
SimpleDDL-MD-Chat 65.0 67.4 71.6 54.7
SimpleDDL-MD-Complete 63.3 61.7 66.0 50.2
@@
No. of GT Tables ChatGPT SQLCoder-34B InternLM-70B Codellama-34B Llama2-Chat-70B InternLM2-20B
1 62.00 44.00 44.00 32.00 22.00 40.00
@@
Figure 3: Word cloud representation of error information for incorrect SQL queries generated by different LLMs. Top: System Error and Result Error. Bottom: Detailed classification of Result Error.

### v2

@@ Abstract
Large Language Models (LLMs) have emerged as a powerful tool in advancing the Text-to-SQL task, significantly outperforming traditional methods. Nevertheless, as a nascent research field, there is still no consensus on the optimal prompt templates and design frameworks. Additionally, existing benchmarks inadequately explore the performance of LLMs across the various sub-tasks of the Text-to-SQL process, which hinders the assessment of LLMs’ cognitive capabilities and the optimization of LLM-based solutions. To address the aforementioned issues, we firstly construct a new dataset designed to mitigate the risk of overfitting in LLMs. Then we formulate five evaluation tasks to comprehensively assess the performance of diverse methods across various LLMs throughout the Text-to-SQL process. Our study highlights the performance disparities among LLMs and proposes optimal in-context learning solutions tailored to each task. These findings offer valuable insights for facilitating the development of LLM-based Text-to-SQL systems.
@@ 1 Introduction
Text-to-SQL, which involves the automatic transformation of natural language (NL) questions into structured SQL statements, is a pivotal component in facilitating seamless user interaction with databases Qin et al. (2022). Previous approaches to this task primarily focus on pattern matching between natural language and SQL statements, utilizing machine learning models to acquire the mapping between the two Zhong et al. (2017); Li et al. (2023a). However, the introduction and rapid advancement of Large Language Models (LLMs) have brought about a substantial transformation in this field Dong et al. (2023); Pourreza and Rafiei (2024). LLMs have emerged as powerful tools Ruan et al. (2023); Kong et al. (2023); Sui et al. (2023), showcasing tremendous potential in comprehending complex NL questions and generating accurate SQL statements. By combining advanced reasoning techniques and in-context learning capabilities, LLMs have significantly pushed the boundaries of the state-of-the-art in this domain, outperforming traditional methods by a considerable margin.
Despite the continuous improvement of LLM-based methods in various benchmarks such as Spider Yu et al. (2018) and BIRD Li et al. (2024), there remains a critical gap in the systematic benchmarking of these solutions Katsogiannis-Meimarakis and Koutrika (2023); Qin et al. (2022); Kumar et al. (2022). The absence of a comprehensive framework for evaluating LLMs in Text-to-SQL complicates the design and assessment of effective systems. The risk of overfitting, particularly for LLMs trained on coding tasks and open-source datasets, poses a significant challenge to the reliability of benchmark evaluations Roziere et al. (2023). Additionally, the optimal prompt engineering strategies that play a crucial role in guiding LLMs to generate accurate SQL queries are yet to be determined. Although various design schemes are explored in different methods Gao et al. (2023), there is no consensus on the most effective prompt template. Furthermore, the current benchmarks, while comprehensive in their assessment of end-to-end Text-to-SQL tasks, have not yet provided a detailed exploration of the models’ performance across the various sub-tasks and components of the Text-to-SQL process Pourreza and Rafiei (2023); Liu et al. (2023). A detailed exploration of these sub-tasks is crucial for a thorough evaluation of LLMs’ cognitive capabilities and their role in facilitating the Text-to-SQL process. Therefore, it is necessary to develop a more granular benchmarking approach that can accurately reflect the multifaceted nature of Text-to-SQL and inform the creation of more effective LLM-based solutions.
@@ 2.1 Traditional Learning-based Text-to-SQL Methods
Numerous traditional learning-based Text-to-SQL methods existed before the emergence of LLMs Zhong et al. (2017); Xu et al. (2017); Lyu et al. (2020). These methods can be essentially divided into non-seq2seq and seq2seq methods according to their network architecture. In non-seq2seq methods, representative works Wang et al. (2019); Cao et al. (2021); Cai et al. (2021); Hui et al. (2022) typically employ a relation-aware self-attention mechanism as encoder to learn representations of questions and schemas, and then use a grammar-based decoder to generate the SQL as an abstract syntax tree Yin and Neubig (2017), or utilize a sketch-based decoder to obtain the SQL via slot-filling Dong and Lapata (2018); Hwang et al. (2019); Choi et al. (2021). These methods can further benefit from leveraging pre-trained language models like BERT and its extensions Devlin et al. (2018); Liu et al. (2019); Yin et al. (2020) for input embedding initialization. As another line of research, seq2seq methods Raffel et al. (2020); Lin et al. (2020); Qi et al. (2022); Li et al. (2023c) directly translate NL questions into SQL queries through transformer-based Vaswani et al. (2017) architectures in an end-to-end manner. These methods obtain competitive performance by fine-tuning rather than training from scratch, with less effort. They typically use smarter decoding techniques (e.g., constraining the predictions of the decoder Scholak et al. (2021), or schema-aware denoising Xu et al. (2021)) to prevent the production of invalid SQLs. RESDSQL Li et al. (2023a) further decouples the intertwined process of schema linking (determining the schema items like tables and columns in a SQL) and skeleton parsing (determining the SQL keywords) , which alleviates the difficulty of Text-to-SQL. Apart from network architectures, the paradigms employed by these methods vary greatly in terms of input encoding, output decoding, neural training, output refinement, making Text-to-SQL a flourishing research area Katsogiannis-Meimarakis and Koutrika (2023). However, the highest accuracy on the Spider leaderboard achieved by traditional learning-based methods is 79.9%, which is still far from being a reliable Text-to-SQL parser.
@@ 2.2 LLM-based Text-to-SQL Methods
With the advancements in LLMs, researchers are increasingly interest in creating a natural language interface for relational databases through the powerful linguistic and coding capabilities of LLMs, forge a new trend of LLM-based Text-to-SQL. Given the powerful zero-shot reasoning and domain generalization abilities of LLMs, these methods successively refresh the record on the cross-domain Spider leaderboard. C3 Dong et al. (2023), a zero-shot Text-to-SQL method built upon ChatGPT, provides treatment in terms of model input, bias and output, reaching an 82.3% execution accuracy on Spider leaderboard. DIN-SQL Pourreza and Rafiei (2024) proposes decomposing the Text-to-SQL task into smaller sub-tasks effectively and achieves an accuracy of 85.3% on Spider. DAIL-SQL Gao et al. (2023) refreshes the accuracy of Spider with 86.6% through both supervised fine-tuning and a systematic study of in-context learning. It explores how to select the most helpful example and organize them properly in prompts in a few-shot scenario. Similarly, other studies investigate the selection of few-shot demonstrations by synthesising in-domain examples Chang and Fosler-Lussier (2023) and retrieving question skeletons Guo et al. (2023). MAC-SQL Wang et al. (2023a) utilizes multi-agent collaboration for Text-to-SQL tasks and reach an accuracy of 59.6% on more challenging BIRD.
Several other studies focus on special yet non-trivial scenarios, aiming to expand the scope of the Text-to-SQL task. DBCopilot Wang et al. (2023b) considers a more realistic problem, wherein it deals with large-scale schemas, characterized by massive databases and a large number of tables. It proposes the use of a lightweight seq2seq copilot model for schema-routing, increasing scalability in comparison to traditional schema-linking. ACT-SQL Zhang et al. (2023b) designs a chain-of-thought (CoT) Wei et al. (2022) prompt to improve the reasoning ability when generating SQLs, and extended to the multi-turn Text-to-SQL task Li et al. (2023b).
In summary, these methods primarily focus on improving the performance of the overall Text-to-SQL task through several aspects of the sub-tasks of interest in this paper. However, these methods 1) do not individually and systematically study the performance of sub-tasks, and 2) typically rely on experiments using only OpenAI LLMs (ChatGPT, GPT4, etc.), without verifying the robustness of their approaches when apply to other open-source LLMs.
@@ 2.3 Datasets and Evaluation Metrics
WikiSQL Zhong et al. (2017) is considered the first large-scale dataset enabling the training and evaluation of learning-based Text-to-SQL methods, as well as offering a standardized benchmark for straightforward comparison across various methods. It is also known as a cross-domain dataset, featuring over 25,000 tables and 80,000 question-SQL pairs comprising various domains derived from Wikipedia. However, the SQL queries in WikiSQL exhibit low complexity.
@@
Subsequently, most of the recent Text-to-SQL works are done on Spider Yu et al. (2018) because it is widely acknowledged as the most challenging cross-domain benchmark. It comprises 10,181 queries covering 138 different domains, involves multi-table queries (embodied by JOIN), complex SQL clauses (ORDER BY, GROUP BY, and HAVING, etc.) and nested SQLs. Several variants of Spider have been designed to evaluate the adaptability of Text-to-SQL methods Shaw et al. (2020); Gan et al. (2022). Spider-Realistic Deng et al. (2020) removes the explicit mention of column names in the NL questions while keeping the SQL queries unchanged, which is more aligned with the use-case of Text-to-SQL in practice. Spider-Syn Gan et al. (2021a) replaces some schema words in the NL question with their synonyms that reflect real-world question paraphrases. Eliminating such explicit correspondence between NL questions and table schemas poses a greater challenge of Text-to-SQL methods. Spider-DK Gan et al. (2021b) integrates artificial domain knowledge, aiming to investigate the robustness of text-to-SQL models when the questions require rarely observed domain knowledge.
@@
Given that Spider is specifically designed for benchmarking Text-to-SQL methods while diverging from practical scenarios, KaggleDBQA Lee et al. (2021) constructs a small-scale cross-domain dataset from Kaggle, highlighting the diversity of actual web-sourced databases. It comprises knowledge such as documentation and metadata of databases, raising an question of how this extra information could be used to improve the performance.
@@
The newest seminal benchmark is BIRD Li et al. (2024), involving 12,751 Text-to-SQL pairs and 95 databases with a size of 33.4 GB. It incorporates the advantages of previous datasets, such as Spider (which is cross-domain with complex SQLs) and KaggleDBQA (which requires the ability to use external knowledge evidence). It is the first to be curated for evaluating SQL execution efficiency in large-scale databases, thereby further bridging the gap between academic setups and real-world applications. In this paper, we construct a novel dataset built upon BIRD and use it in the evaluation (see Section 3.3).
@@
Two primary evaluation metrics for assessing the accuracy of SQLs on Spider are Exact Matching(EM) and Execution Accuracy(EX). EM measures whether the predicted query as a whole is equivalent to the gold query. It is possible to encounter false negative evaluations since a question might be solvable by multiple syntactically different but semantically identical SQL statements. EX is a more widely used metric, measures whether the result of executing the predicted query matches the gold value. We use EX to evaluate the accuracy of SQLs in this paper. Additionally, BIRD further proposes Valid Efficiency Score (VES), an integrated metric assessing both accuracy of execution results (i.e., EX) and the execution efficiency of SQL queries. To increase the VES, methods are required to enhance both execution accuracy and efficiency of SQL queries. We use this metric to assess the SQL optimization in Section 4.3. Refer to Appendix C.1 for detailed definitions of evaluation metrics.
@@ 3.2 Evaluation Models
Our benchmarking study evaluates the performance of two distinct categories of LLMs with varying parameter sizes: general-purpose and coding-specific. General-purpose LLMs are designed for versatile text generation and comprehension across diverse domains, trained on extensive internet text datasets. Specifically, ChatGPT (gpt-35-turbo-16k)  [^1] https://chat.openai.com, LLaMa2-Chat-70B Touvron et al. (2023), InternLM-70B  [^2] https://internlm.intern-ai.org.cnand InternLM2-20B are selected as the main baseline models. Coding-specific LLMs are fine-tuned and optimized for programming scenarios, excelling in code generation and technical language understanding. In this paper, the performance analysis of Codellama-34B Roziere et al. (2023) and SQLCoder-34B  [^3] https://github.com/defog-ai/sqlcoder are provided.
@@ 3.3 Dataset Construction
Specifically, our construction process involves a systematic analysis of the original BIRD dataset, identifying queries of varying difficulty levels and involving different numbers of tables (1, 2, 3, and more than 3). We modify and expand these queries by altering table and column names, as well as filtering conditions, to create a more diverse set of challenges. In cases where the original dataset lacks sufficient examples with four or more tables (there are only 20 instances in BIRD-Dev dataset), queries that involved three tables are expanded to four. As shown in Table 2, this process generates 50 new instances for each category, resulting in the “BigTable-0.2k” dataset. Moreover, each item in the dataset underwent mutual verification by at least two individuals to ensure the accuracy.
@@ 4.1 Text2SQL
Unlike previous learning-based studies, the primary challenge in LLM-based Text-to-SQL is the design of an effective prompt template $\mathcal{P}$ (as introduced in Section 3.1) for LLMs to generate accurate SQL queries, known as prompt engineering. Researchers have evaluated a variety of prompt templates Gao et al. (2023). However, these representations lack uniformity in their structure, making it difficult to find out how a specific feature within a prompt template impact performance.
@@
prompt template SQLCoder-34B Codellama-34B InternLM-70B Llama2-Chat-70B
DDL-HTML-Chat 57.8 63.7 65.0 49.6
DDL-HTML-Complete 61.8 65.2 53.8 50.2
DDL-MD-Chat 63.2 68.4 66.3 48.7
DDL-MD-Complete 62.4 69.8 64.1 46.8
DDL-Coding-Chat 60.3 67.1 66.1 48.4
DDL-Coding-Complete 59.7 66.9 62.9 53.4
SimpleDDL-MD-Chat 65.0 71.6 67.4 54.7
SimpleDDL-MD-Complete 63.3 66.0 61.7 50.2
@@
No. of GT Tables ChatGPT SQLCoder-34B Codellama-34B InternLM-70B Llama2-Chat-70B InternLM2-20B
1 60.00 44.00 32.00 44.00 22.00 40.00
@@
Figure 3: Word cloud representation of error information for incorrect SQL queries generated by LLMs. Top: System Error and Result Error. Bottom: Detailed classification of Result Error.